# Boosting Text-to-Chart Retrieval with Semantic Insights

### v2

@@
Abstract.
Charts are crucial for data analysis and decision-making. Text-to-chart retrieval systems have become increasingly important for Business Intelligence (BI), where users need to find relevant charts that match their analytical needs. These needs can be categorized into precise queries that are well-specified and fuzzy queries that are more exploratory – both require understanding the semantics and context of the charts. However, existing text-to-chart retrieval solutions often fail to capture the semantic content and contextual information of charts, primarily due to the lack of comprehensive metadata (or semantic insights). To address this limitation, we propose a training data development pipeline that automatically synthesizes hierarchical semantic insights for charts, covering visual patterns (visual-oriented), statistical properties (statistics-oriented), and practical applications (task-oriented), which produces 207,498 semantic insights for 69,166 charts. Based on these, we train a CLIP-based model named ChartFinder to learn better representations of charts for text-to-chart retrieval. Our method leverages rich semantic insights during the training phase to develop a model that understands both visual and semantic aspects of charts. To evaluate text-to-chart retrieval performance, we curate the first benchmark, CRBench, for this task with 21,862 charts and 326 text queries from real-world BI applications, with ground-truth labels verified by the crowd workers. Experiments show that ChartFinder significantly outperforms existing methods in text-to-chart retrieval tasks across various settings. For precise queries, ChartFinder achieves up to 66.9% NDCG@10, which is 11.58% higher than state-of-the-art models. In fuzzy query tasks, our method also demonstrates consistent improvements, with an average increase of 5% across nearly all metrics.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Visualization charts (charts for short) are essential for various fields, including data analysis and business decision-making (Wu et al., 2024; Luo et al., 2018a; Shen et al., 2022; Qin et al., 2020). In particular, Business Intelligence (BI) systems heavily rely on charts to present complex data in ways that provide users with intuitive and actionable insights (Setlur et al., 2023; Luo et al., 2023). Most charts are carefully designed to serve specific analytical tasks, with considerations for visual design, chart types, and data relationships.
Given their quality and versatility, these well-crafted charts are valuable resources that can be reused across different tasks and contexts. Given a user-provided text query and a chart repository, the text-to-chart retrieval task is to find a set of charts that align with the user’s query. This process requires understanding both the visual characteristics of the charts and the semantic content of the user query, ensuring that the retrieved charts are not only relevant but also contextually suitable for the user’s analytical task.
Figure 1. ChartFinder integrates both visual and semantic aspects of charts to improve text-to-chart retrieval.
Text-to-Chart Retrieval in Real-world BI Systems. Through collaboration with a leading BI company $X$ [Redacted], we analyzed real-world user behavior and identified two primary categories of user queries for chart retrieval: precise queries and fuzzy queries. These categories reflect the diverse needs of users, ranging from specific, data-driven queries to more exploratory or stylistic tasks.
Precise queries are characterized by data-driven intent and focus on specific analytical targets. These queries aim to retrieve charts that present well-defined information, such as trends, comparisons, or aggregated data points, to answer a clear analytical question. For example, as shown in Figure 1, a user may ask, “Find charts showing the hotel booking distribution across channels in Europe for the year 2022.” Such queries demand high semantic alignment to ensure the retrieved charts align closely with the user’s specific requirements. To achieve this, retrieval systems must effectively leverage both the data context and the visual features of chart images to establish strong semantic relationships between the query and the chart.
Fuzzy queries, on the other hand, are more open-ended and often focus on exploratory tasks and abstract information. These queries are less constrained and may emphasize stylistic or structural elements of charts rather than specific data content. For example, a user might query, “Use a pie chart to compare different sale methods”, as shown in Figure 1. Addressing such queries requires a retrieval system capable of understanding a broader intent and contextual flexibility inherent in exploratory tasks.
Limitations of Existing Methods. Text-to-image retrieval has seen remarkable advancements, leveraging state-of-the-art Multimodal Large Language Models (MLLMs) (Liu et al., 2022). These approaches have demonstrated effectiveness in general image retrieval tasks by efficiently capturing visual features from images. However, when being directly adapted to text-to-chart retrieval, these methods face several challenges in addressing the unique aspects of charts.
First, existing methods fail to capture the semantic depth of charts. Unlike natural images, which primarily convey visual information, charts encode complex data relationships, trends, and comparisons. Existing methods that focus solely on visual features often retrieve charts that are visually similar but semantically irrelevant. For example, two charts may share visual characteristics (e.g., both being bar charts) but represent entirely different datasets, topics, or analytical purposes. This limitation is particularly problematic for precise queries, which require high semantic alignment between the query and the retrieved chart to meet data-driven needs.
Second, current approaches fail to address both precise and fuzzy queries effectively. For precise queries, existing methods that prioritize visual features lack the ability to connect queries with the underlying data or analytical intent, leading to poor retrieval performance. For fuzzy queries, existing methods struggle to interpret broader user intents, such as stylistic preferences or exploratory goals, as they lack mechanisms to capture design elements of charts.
Key Idea. To improve text-to-chart retrieval, it is essential to capture both the visual and semantic aspects of charts. Our key idea is to extract semantic insights from chart metadata (e.g., source code) and combine this with the chart images. By training a model that incorporates both the charts (in the form of images) and the semantic insights (in the form of text), we aim to develop a system that comprehensively understands and bridges the gap between visual representation and the underlying semantic meaning of charts.
However, implementing this idea presents several challenges.
First, manually annotating semantic insights for charts is costly and time-consuming. Thus, we need to develop an efficient and scalable method to derive these semantic insights from chart metadata automatically. Once semantic insights are generated, the second challenge is to effectively integrate them with the visual features of the charts. We need an effective model architecture and training strategy that can seamlessly combine these two modalities, i.e., semantic insights and visual features, to generate more accurate representations of the charts. Third, there is currently no readily available corpus containing text queries and their corresponding charts to evaluate the performance of text-to-chart retrieval.
Our Proposal. We propose ChartFinder, a new text-to-chart retrieval model that leverages contrastive learning techniques to align both the visual features and semantic insights of charts, based on the Contrastive Language-Image Pre-Training architecture (Radford et al., 2021; Zhang et al., 2025). Specifically, our model is trained with both chart images (visual content) and their corresponding semantic insights (textual content), as shown in Figure 1. It is important to note that ChartFinder does not require synthesized semantic insights during the retrieval phase. Instead, it relies solely on text queries and candidate charts to perform the retrieval. The semantic insights are utilized exclusively during the training phase, where they help the model better align chart images with their semantic meaning, thus enhancing its ability to retrieve relevant charts from a repository based on textual descriptions.
Once trained, ChartFinder accepts a user query and a chart repository as input, retrieving the top-$k$ most relevant charts by calculating the embedding distance between the text query and the charts.
To train the model effectively, we have developed a training data generation pipeline that automatically generates a large number of charts along with their corresponding semantic insights. Inspired by techniques in automatic data visualization (Luo et al., 2018a) and insight generation (Ma et al., 2021), which create insightful charts from raw data, our pipeline synthesizes three key levels of semantic insights: (1) Visual-oriented Insight: a high-level summary of the chart, capturing the overall pattern and visual trends; (2) Statistics-oriented Insights: an in-depth analysis of the trends, comparisons, and relationships within the data, highlighting statistical patterns and anomalies; and (3) Task-oriented Insights: contextual information on how the chart can be applied for suitable data analysis tasks.
@@ 1. Introduction
To evaluate the performance of the text-to-chart retrieval task, we introduce CRBench, the first text-to-chart retrieval benchmark sourced from real-world BI scenarios. CRBench contains 326 queries and 21,862 chart images. All query-chart labels are verified by crowd workers in order to capture the diversity of user needs.
Contributions. We make the following contributions:
@@
(2) ChartFinder Model. We propose ChartFinder, a novel text-to-chart retrieval model that combines contrastive learning with the CLIP architecture to align visual features and semantic insights of charts. The model leverages the synthesized training data to improve retrieval accuracy, making it more context-aware in handling both precise and fuzzy queries (Section 4).
(3) CRBench: A New Benchmark. We curate CRBench, the first benchmark for evaluating text-to-chart retrieval sourced from real-world BI scenarios. It includes 326 text queries and 21,862 charts, with query-chart labels verified by crowd workers to capture the diversity of real user needs (Section 5).
@@
(4) Extensive Experiments. Our experiments show that the synthesized training data pipeline significantly enhances the performance of ChartFinder, outperforming existing methods across multiple benchmarks. This includes demonstrating improved performance in zero-shot scenarios. The results validate the effectiveness of both the training data synthesis pipeline and the ChartFinder model in real-world text-to-chart retrieval tasks (Section 6).
2. Preliminaries
Charts. A chart $C$ is a visual representation of data in the form of rasterized images designed to communicate patterns, trends, or comparisons across one or more variables.
Chart Metadata. Let $M=(T,X,Y,CT,D,E)$ be the associated metadata of a chart, where:
  • $T$ is the title of the chart,
  • $X$ and $Y$ are the labels for the x-axis and y-axis, respectively,
  • $CT\in\{\text{line},\text{bar},\text{scatter},\text{pie},\dots\}$ is the chart type,
  • $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ denotes the underlying data, where $(x_{i},y_{i})$ are data points, and $n$ is the number of data points,
  • $E$ represents additional elements such as the legend, data labels, or annotations.
Semantic Insights of Charts. As shown in Figure 1 ( [Uncaptioned image] area), the information content of a typical chart can be divided into three key components: (1) the visual representation of the data, (2) the statistical features embedded in the data, and (3) the application or contextual value the chart provides. These components correspond to the visual expression, numerical basis, and purpose of the chart, respectively. To ensure that the model fully understands the chart’s semantics, we have designed three types of semantic insights. Given a chart $C$ and its metadata $M$, we define three levels of semantic insights $I=(I^{v},I^{t},I^{s})$, as follows.
  • $I^{v}$ represents Visual-oriented Insights: These insights are derived from visual patterns, trends, or relationships observed in the chart. For example, trends shown in the chart may reveal increasing or decreasing patterns over time
  • $I^{t}$ denotes Task-oriented Insights: These insights relate to the practical applications of the chart in specific tasks or decision-making processes. For example, a chart illustrating healthcare trends could help policymakers identify areas of resource allocation.
  • $I^{s}$ represents Statistics-oriented Insights: These are quantitative analyses and statistical observations derived from the data. For example, identifying correlations, distributions, or outliers in the dataset used for the chart.
The process of deriving semantic insights is represented as $f:(C,M)\rightarrow I$, where $f$ analyzes the chart $C$ and its metadata $M$ to produce the insights $I$. Please refer to Section 3 for details.
Example 0.
(Example of Chart, Metadata, and Semantic Insights.) As shown in Figure 2, the line chart $C$ is rendered using the Plotly visualization library. The metadata for this chart is the source code used by Plotly, which contains the title, axis labels, data points, chart type, and other relevant information. We utilize the LLaMA-3.1 model to analyze this metadata and generate semantic insights.
Text-to-Chart Retrieval. Let $\mathcal{C}=\{C_{1},C_{2},\dots,C_{n}\}$ denote a repository of charts and a textual query $Q$ provided by a user. The objective of the text-based chart retrieval problem is to retrieve a ranked list of $k$ charts, $\mathcal{C}_{k}=\{C_{r_{1}},C_{r_{2}},\dots,C_{r_{k}}\}\subseteq\mathcal{C}$, where $k\ll n$ and $C_{r_{i}}$ represents the $i$-th most relevant chart to the query $Q$.
3. Semantic Insights Synthesis Pipeline
Figure 2. An overview of the semantic insights synthesis pipeline.

### v4

# Boosting Text-to-Chart Retrieval through Training with Synthesized Semantic Insights

@@
Abstract
Text-to-chart retrieval, enabling users to find relevant charts via natural language queries, has gained significant attention. However, evaluating models in real-world business intelligence (BI) scenarios is challenging, as current benchmarks fail to simulate realistic user queries or test for deep semantic understanding with static chart images.To address this gap, we introduce CRBench, the first real-world BI-sourced benchmark comprising 21,862 charts and 326 queries, utilizing a Target-and-Distractor paradigm to evaluate discriminative retrieval among highly similar candidates. Testing on CRBench reveals that existing methods, which rely primarily on visual features, perform poorly and fail to capture the rich analytical semantics of charts. To address this performance bottleneck, we propose a semantic insights synthesis pipeline that automatically generates three hierarchical levels of insights for charts: visual patterns, statistical properties, and practical applications. Using this pipeline, we produced 207,498 semantic insights for 69,166 charts as training data. By leveraging this data to bridge the gap between natural language query intent and latent visual representations via multi-level semantic supervision, we develop ChartFinder, a specialized model capable of deep cross-model reasoning. Experimental results show ChartFinder significantly outperforms state-of-the-art methods on CRBench, achieving up to 66.9% NDCG@10 for precise queries (an 11.58% improvement) and an average increase of 5% across nearly all metrics for fuzzy queries. This work provides the community with a much-needed benchmark for realistic evaluation and demonstrates a powerful data synthesis paradigm for enhancing a model’s semantic understanding of charts.
Boosting Text-to-Chart Retrieval through Training with Synthesized Semantic Insights
Yifan Wu1  Lutao Yan1  Yizhang Zhu1  Yenchi Tseng1 Yinan Mei2 Yong Wang2 Jiannan Wang2,3 Nan Tang1 Yuyu Luo1thanks: Yuyu Luo is the corresponding author. yuyuluo@hkust-gz.edu.cn 1HKUST(GZ), China  2Huawei Cloud BU, China  3Simon Fraser University, Canada
1 Introduction
Visualization charts are crucial tools in data analysis and business decision-making, providing intuitive ways for users to understand complex data Wu et al. (2024); Luo et al. (2018a); Shen et al. (2022); Qin et al. (2020). Consequently, text-to-chart retrieval systems that allow users to find relevant charts based on natural language queries (i.e., text queries) have become increasingly valuable, attracting significant attention from both the database community Ji et al. (2024b, a); Luo et al. (2023) and industry practitioners Setlur et al. (2023). In real-world BI scenarios, queries range from precise queries targeting specific analytical information (e.g., “Find charts showing hotel booking distribution in Europe for 2022”) to fuzzy queries for exploratory analysis (e.g., “Use a pie chart to compare different sale methods”) Ji et al. (2024a, b). Both require a deep understanding of the rich analytical semantics embedded in charts, such as data relationships and statistical trends Qin et al. (2020). This challenge is exacerbated when charts exist only as static images without accessible metadata, a common scenario where visual-based methods inherently struggle Luo et al. (2023); Ji et al. (2024a); Li et al. (2022a).
A core obstacle impeding progress in this field is the lack of a dedicated benchmark for realistic evaluation. Existing chart-related datasets are primarily designed for tasks like captioning and fail to simulate real-world retrieval challenges. They lack concise, intent-driven queries and a large repository of visually and semantically similar charts to test a model’s true discriminative power. To address this critical evaluation gap, we first curate and introduce CRBench, the first benchmark for text-to-chart retrieval sourced from real-world BI scenarios. CRBench contains 326 text queries and 21,862 chart images, with all query-chart labels verified by crowd workers to capture the diversity of real user needs.
Benchmarking existing models, including powerful Multimodal Large Language Models (MLLMs) Liu et al. (2022); Xiao et al. (2024b); Zhou et al. (2024b); Zhang et al. (2024), on CRBench reveals their significant limitations. These models struggle to capture the deep analytical semantics required to differentiate between challenging charts. To address this performance bottleneck, we argue that the key lies in providing higher-quality training data that explicitly teaches models to bridge the gap between visual features and their analytical meaning. To this end, we propose our second major contribution: an automatic semantic insights synthesis pipeline. Inspired by automatic data visualization Luo et al. (2018a) and insight generation methods Ma et al. (2021), this pipeline automatically synthesizes three hierarchical levels of insights: (1) Visual-oriented, (2) Statistics-oriented, and (3) Task-oriented.
To validate the challenge of our benchmark and the effectiveness of our data pipeline, we use our synthesized data to fine-tune several CLIP-based models. We name the model resulting from fine-tuning Long-CLIP Zhang et al. (2025) as ChartFinder. The core idea is to leverage the synthesized insights exclusively during the training stage to learn robust visual-semantic representations. During retrieval, the model relies solely on chart images.
Contributions. We make the following contributions.
@@ 1 Introduction
(1) CRBench Benchmark. We curate CRBench, the first benchmark for evaluating text-to-chart retrieval sourced from real-world BI scenarios. It includes 326 text queries and 21,862 charts, with query-chart labels verified by crowd workers to capture the diversity of real user needs (Section 3).
@@
(3) A High-Performance Model Validating the Pipeline. By fine-tuning the CLIP architecture with our synthesized data, we produce ChartFinder. It not only validates the effectiveness of our data synthesis pipeline by establishing a new state-of-the-art but also serves as a strong baseline for future research on CRBench.
@@
(4) Extensive Experiments. Extensive experiments validate both the challenge of CRBench as a benchmark and the effectiveness of our synthesis pipeline. The results demonstrate that our synthesized data significantly improves the performance of multiple CLIP-based architectures across several benchmarks, including in challenging zero-shot scenarios (Section 5).
2 Related Work
Figure 1: The construction workflow of the CRBench.
Chart Retrieval. Chart retrieval aims to find similar charts from a repository using a chart as a query. Existing methods follow two main paradigms: perception-based and data-based. Perception-based approaches analyze the chart image, using either visual style and structure Saleh et al. (2015); Hoque and Agrawala (2019), deep visual embeddings Luo et al. (2023), or hand-drawn sketches of trendlines to find matches Mannino and Abouzied (2018). In contrast, data-based approaches assume access to the underlying data of the query chart and use it to guide the retrieval process Wang and Palpanas (2021); Lekschas et al. (2020); Siddiqui et al. (2016). Chart search has similarities with text-to-chart retrieval, particularly in terms of underlying semantic complexity. However, our work fundamentally differs by addressing the additional challenge of aligning natural language queries with chart semantics, especially when metadata is unavailable.
Text-to-Chart Retrieval. Previous work on text-to-chart retrieval has largely followed two paths: visualization recommendation systems that guide user exploration Vartak et al. (2017); Ye et al. (2024); Qin et al. (2018b, a); Bendeck et al. (2024); Srinivasan and Setlur (2021) and natural language interfaces that accept analytical queries Shen et al. (2022); Tang et al. (2022); Luo et al. (2018b, 2022b, 2021); Setlur et al. (2023); Srinivasan and Setlur (2023). While significant progress has been made, these methods often struggle to fully understand user intent. We believe embedding queries and charts into a shared latent space offers a more robust solution. The development of Multimodal Large Language Models (MLLMs) Li et al. (2022b, 2023); Yuan et al. (2021); Xiao et al. (2024a) and pre-trained models like CLIP Radford et al. (2021) has enabled text-to-chart retrieval to be treated as a cross-modal task. Numerous variants have been developed to enhance CLIP’s performance, including those with improved visual encoders like EVA-CLIP Sun et al. (2023); Fang et al. (2024), better text encoders like Jina-CLIP-v2 Xiao et al. (2024b); Koukounas et al. (2024); Sturua et al. (2024), or models improved through large-scale fine-tuning such as UniVL-DR Liu et al. (2022), MagicLens Zhang et al. (2024), VISTA Zhou et al. (2024a), and MARVEL Zhou et al. (2024b). However, research focusing MLLM-based methods specifically on chart retrieval remains limited. This paper addresses that gap by using contrastive learning to inject deep semantic insights into our model, ChartFinder.
3 CRBench Benchmark
In this section, we first discuss the motivation for curating a new text-to-chart retrieval benchmark (Section 3.1), followed by an introduction to the benchmark curation process (Section 3.2). After that, we introduce the key characteristics of the first text-to-chart benchmark CRBench (Section 3.3). Finally, we discuss the data quality and limitations of CRBench.
3.1 Motivation and Design Consideration
To rigorously evaluate text-to-chart retrieval models, a specialized benchmark is required. However, existing chart-related datasets are ill-suited for this task. While chart-to-text datasets like VisText Tang et al. (2023) and Chart-to-Text Kantharaj et al. (2022) are valuable resources, they were fundamentally designed for generation tasks, such as captioning, not retrieval. This creates several critical limitations when they are repurposed for retrieval.
First, their “queries” are typically long, descriptive captions, rather than the concise, intent-driven natural language phrases in real-world BI systems. Second, and more importantly, they often feature a one-to-one or one-to-two chart-to-text ratio. This kind of ratio fails to simulate a realistic retrieval environment, where a single user query can be matched against a massive repository containing thousands of visually and semantically similar charts.
Consequently, a significant gap exists for a benchmark that can evaluate a model’s performance on realistic, intent-driven queries against a large and challenging chart repository.
Design Consideration. To fill this gap, we aim to propose a real-world dataset designed for text-to-chart retrieval, enabling the fair comparison of retrieval models across various use cases. Several key factors were considered, as outlined below:
(1) Precise and Fuzzy Queries: As discussed in Section 1, user queries in real-world applications can be both precise (focused on specific details) and fuzzy (more general). Our benchmark includes both types of queries to reflect real-world user behavior and more accurately evaluate model performance across varying query complexities.
(2) Distributional Separation: To ensure a fair and rigorous evaluation of model generalization, we maintain a strict separation between the data sources of our training set and the benchmark itself. Therefore, while our training data is generated from Kaggle datasets, the charts for CRBench are sourced exclusively from different, real-world applications not used in training.
(3) Query Realism and Unambiguity: The benchmark must feature queries that reflect the full spectrum of real-world user behavior, including both precise and fuzzy types. Crucially, to enable rigorous and reliable evaluation, each query must be constructed to have a single, unambiguous ground-truth chart, even in the presence of highly similar distractors.
3.2 Benchmark Construction
The key challenge in constructing CRBench is ensuring each query has a clear, unambiguous ground-truth chart, particularly when evaluating retrieval performance on visually similar charts. To systematically address this, we adopt a clearly defined four-stage methodology, as illustrated in the Figure 1.
Step 1: Real-world Chart Collection. To ensure CRBench reflects real-world chart retrieval scenarios, we collected charts from reliable sources widely used in business intelligence and data analysis: (i) Tableau: A leading data visualization platform for charting and business analytics Tableau (2003). (ii) Statista: A platform providing statistical data across various sectors statista (2007). (iii) Pew Research Center: A non-profit organization focusing on social science research Research (2004). (iv) Our World in Data: An open-access platform researching global challenges through charts on topics like economics, health, and society OWID (2011). From these sources, we gathered 21,862 charts covering various domains, including finance, business, e-commerce, technology, and more.
Step 2: Similarity-based Grouping. To construct a challenging retrieval environment based on the “target-and-distractor” paradigm, we first embedded all charts into 768-dimensional vectors using a pre-trained ViT and computed their pairwise cosine similarities.
The grouping process was then performed as follows: we iterated through each chart and identified its four most visually similar counterparts. A challenging group was formed only if the similarity score between this initial chart and all four of its counterparts was $\geq 0.90$. If this strict condition was met, the initial chart that served as the anchor for the group was designated as the target chart, with the other four serving as distractors. Through this systematic process, we ultimately generated 247 distinct groups, each containing one clearly defined target and four highly similar distractors.
Step 3: Text Query Generation.
To generate high-quality and unambiguous queries, we initiated the process with an exploratory crowdsourcing study on the Appen platform. In this study, we presented crowdworkers with the previously created groups of five visually similar charts (one target and four distractors). Their task was to author a query for the target chart that could uniquely distinguish it from the other four.
However, we found that the quality of these manually generated queries was often poor. Many queries were either too general to be effective for differentiation or were simply literal paraphrases of the chart’s title. This outcome highlighted the need for a more controlled and reliable generation process.
Therefore, we transitioned to a controlled query-generation methodology using GPT-4o. Crucially, the prompt designed for GPT-4o was informed by the patterns and shortcomings observed in the initial human-generated queries from the first step. The model was explicitly instructed to generate precise and fuzzy queries aimed at distinguishing a given target chart from its four distractors, thereby overcoming the generality issues we had identified and ensuring the queries were highly targeted.
This automated process is remarkably cost-effective; generating queries for all 247 groups incurred a total API cost of approximately \$2.81, utilizing roughly 1.09 million tokens.
Step 4: Human Consensus Validation. Finally, we performed a stringent multi-worker validation to ensure each query was unambiguous. Each GPT-4o-generated query, along with its corresponding five similar charts (one target and four distractors), was reviewed by nine independent crowdworkers. A query was accepted into CRBench only if a strong consensus of at least five out of the nine workers was reached. Through this rigorous quality control, we ultimately retained 195 precise queries and 131 fuzzy queries, each validated to have a single, unambiguous ground-truth chart.
Table 1: Comparison of Existing Benchmarks. The statistics for Chart-to-Text and VisText are from the respective test sets. The column “C/Q” indicates the ratio for charts (C) and text queries (Q). The column “Dist. of Query Len.” Shows the distribution of lengths of text queries in characters. The column “CT” means the types of charts.
Datasets CT #-Charts #-Queries C/Q Dist. of Query Len. Samples of Similar Charts
VisText 3 882 1764 0.5 [Uncaptioned image] [Uncaptioned image]
Chart-to-Text 6 1393 1393 1 [Uncaptioned image] [Uncaptioned image]
CRBench 11 21,862 326 67 [Uncaptioned image] [Uncaptioned image]
3.3 CRBench Overview
In this section, we compare CRBench with two existing benchmarks, Chart-to-Text Kantharaj et al. (2022) and VisText Tang et al. (2023), both of which focus on chart-related tasks like caption generation and description. While these benchmarks are useful for evaluating chart captioning and description generation, they do not address the text-to-chart retrieval task. In our study, we repurpose the captions in these benchmarks as queries for retrieval tasks, enabling us to compare their effectiveness with CRBench in retrieving relevant charts based on text queries. Specifically, in VisText, each chart is paired with two types of captions: L1 captions (e.g., basic properties) and L2+L3 captions (e.g., trends and statistics). This results in a chart-to-query ratio of 1:2. In contrast, Chart-to-Text pairs each chart with a single caption, resulting in a 1:1 ratio. In CRBench, however, the ratio is much higher (67:1), meaning the number of charts significantly exceeds the number of queries, which is more representative of real-world scenarios where charts vastly outnumber queries.
Moreover, we compared examples of similar charts across the benchmarks. In VisText, similar charts show strong visual resemblance but often differ significantly in semantic content. In Chart-to-Text, charts deemed similar tend to align well semantically but differ visually. On the other hand, CRBench ensures that similar charts are closely aligned both visually and semantically, posing a significant challenge for models to retrieve the ground truth. Overall, CRBench presents a more challenging and realistic benchmark for evaluating text-to-chart retrieval models.
3.4 Data Quality and Limitations
While crowdsourcing was used to generate queries, we addressed quality issues through a second round of crowdsourcing and GPT-4o. This process helped ensure the relevance and clarity of the queries, resulting in a high-quality set of precise and fuzzy queries.
Despite careful curation, crowdsourcing variability and the evolving landscape of data visualizations may require periodic updates to CRBench to maintain its relevance and accuracy.
4 Semantic Insights Synthesis Pipeline
Figure 2: An overview of the semantic insights synthesis pipeline.